OpenStudy (anonymous):

Simplify: √98x a. 7x√2 b. 2x√7 c. 7√2x d. 2√7x

OpenStudy (anonymous):

c

OpenStudy (anonymous):

√98x = √49*√2x = 7√2x

OpenStudy (anonymous):

separate it into 2 square roots, one with perfect squares and one imperfect \[\sqrt{49}\sqrt{2x}\]

OpenStudy (anonymous):

now simplify the first term \[7\sqrt{2x}\]

OpenStudy (anonymous):

voila